A sheet of gold weighing 9.6 g and at a temperature of 19.2C is placed on a flat sheet of iron weighing 18.1 g and at a temperature of 57.1C. What is the final temperature of the combined metals? Assume no heat was lost to the suroundings
1
Expert's answer
2017-07-12T13:23:06-0400
Q(Au)=-Q(Fe)
m(Au)*C(Au)*(t2-t1(AuS)m(Fe)*C(Fe)*(t2-t1(FeS)
0.096kg*120J⁰C/kg*(t2-19.2)=-0/181kg*460J⁰C/kg*(t2-57.1)
11.52*(t2-19.2)=-83.26(t2-57.1)
11.52t2-221.184=-83.26t2+4754.146
94.78t2=4975.33
t2=52.59⁰C